OpenStudy (anonymous):

factor the expression 3x^3 + 3x^2 + x + 1

OpenStudy (ajprincess):

can u factorise 3x^3 + 3x^2?

OpenStudy (anonymous):

so do 3 x 3 x 3 =27 3 x 2 = 9 and add them together? 36?

OpenStudy (ajprincess):

nope. 3x^3 =3*x*x*x 3x^2=3*x*x

OpenStudy (anonymous):

so 81 and 27? add them?

OpenStudy (ajprincess):

3x^3=3*x*x*x=3x^2*x 3x^2=1*3x^2 so vat is common to both 3x^3 and 3x^2?

OpenStudy (anonymous):

x?

OpenStudy (ajprincess):

3x^2

OpenStudy (ajprincess):